The percentage of wins for a basketball team going into the playoffs was 90% last season.


a. What is the probability that the team sweeps (4-0) the initial best-of-7 playoff series?


b. What is the probability that the team wins the initial best-of-7 playoff series?



1
Expert's answer
2021-11-10T15:50:16-0500

p=0.90

q=0.10

n=7

a. "0.90 \times 0.90 \times 0.90 \times 0.90 = 0.6561"

The probability that the team sweeps (4-0) the initial best-of-7 playoff series is 0.6561

b.

"P(X=x) = C^7_x \times 0.90^x \times 0.1^{7-x} \\

P(wins) = P(X≥4) = P(X=4) + P(X=5) +P(X=6) +P(X=7) \\

= C^7_4 \times 0.90^4 \times 0.1^{7-4} + 

+ C^7_5 \times 0.90^5 \times 0.1^{7-5 } + 

+ C^7_6 \times 0.90^6 \times 0.1^{7-6} + 

+ C^7_7 \times 0.90^7 \times 0.1^{7-7} \\

= 0.0229635 + 0.1240029 + 0.3720087 + 0.4782969 \\

= 0.997272"

The probability that the team wins the initial best-of-7 playoff series is 0.997272